# Hyper-bishops, Hyper-rooks, and Hyper-queens: Percentage of Safe Squares on Higher Dimensional Chess Boards

### v3

@@ Abstract.
Authors were supported by Emmanuel College Cambridge, The William & Mary Charles Center, The Cissy Patterson Fund, Williams College, The Finnerty Fund, and NSF Grant DMS2241623. Special thanks to Warren Johnson for helpful comments on an earlier draft and to the referees for improving the clarity of the writing.
@@ 1. Introduction
Miller et. al: Rooks in 2 Dimensions Section 2: Combinatorial Generalization Section 3: Bishops and Queens in 2 Dimensions Section 4: Line-pieces in Higher Dimensions Section 5: Hyper-pieces in Higher Dimensions
@@
and
@@ 1.1. Notation and Basic Definitions in Higher Dimensions
We provide two natural extensions of chess pieces in higher dimensions, namely line-pieces and hyper-pieces. Line-pieces attack along lines; for example, a line-rook attacks any piece that shares (k1)(k-1) coordinates with it. Hyper-pieces attack along (k1)(k-1)-dimensional hyperplanes; for example, a hyper-rook attacks any piece that shares one coordinate with it.
When considering line-pieces, we determine that the expected proportion of safe squares when randomly placing $n^{k-1}$ on a $k$-dimensional chessboard is $1/e^{k}$. We then consider line-bishops and line-queens and provide an integral for estimating the expected proportion of safe squares as $n$ approaches infinity. We computationally determine the expected proportion of safe squares when $k=3$. Our methods here are similar to the 2-dimensional case.
Finally, we consider hyper-pieces in $k$ dimensions, first considering hyper-rooks, where we prove the following theorem.
@@
As $n$ approaches infinity, the mean number of safe spaces on a $k$-dimensional chessboard with side length $n$ and $n$ randomly placed hyper-rooks is asymptotically $n^{k}/e^{k}$, and the expected proportion of safe spaces converges to $1/e^{k}$.
While a rook at any given space attacks the same number of spaces, the number of spaces a bishop attacks varies widely, often with little symmetry. Due to this, we are unable to determine the expected proportion of safe spaces in $k$ dimensions, and instead analyze $3$ dimensions.
@@ 3.1. Probability of a Safe Square for Bishops and Queens
Interpreting this as a Riemann sum, with $x=r/n$, we can rewrite this as an integral, and then compute via integration by parts.
@@
$=8e^{-2}(\frac{e}{4}-(\frac{e}{4}-\frac{1}{4}))$
$=\frac{2}{e^{2}}$
Because we showed that constant terms do not impact the final limit, we have that the expected proportion of safe spaces is
@@
We can define the even case similarly. The 0th ring has $4$ spaces, and the $r$th ring has $4(2r+1)$ spaces. Because we index from 0, the outermost ring is the (n/21)(n/2-1)st ring. A bishop in the 0th ring attacks n1n-1 spaces along one diagonal and n2n-2 along the other. For every ring movement outward, it attacks $2$ less spaces.
Thus, the expected percentage of safe spaces on an even $n\times n$ chessboard with $n$ randomly placed bishops is
@@
Following the proof above, we evaluate this limit of this expression as $n$ approaches infinity to be $2/e^{4}$.
@@ 3.2. Variance for 2D Bishops and Queens
By the definition of variance and standard properties:
@@
$=\ \frac{1}{n^{2k}}\left(\sum_{i_{1},i_{2},...,i_{k}=1}^{n}\operatorname{{Var}}(X_{i_{1},i_{2},...,i_{k}})+\sum_{i_{1},...,i_{k},j_{1},...,j_{k}=1}^{n}\operatorname{{Cov}}(X_{i_{1},i_{2},...,i_{k}},X_{j_{1},j_{2},...,j_{k}})\right).$ (3.8)
@@
Covariance Term We now consider the covariance term. $X_{i_{1},...,i_{k}}$ and $X_{j_{1},...,j_{k}}$ both consider all possible board configurations, so have the same expected value. Then, the following holds:
@@
$=\ \operatorname{\mathbb{E}}[X_{i_{1},...,i_{k}}X_{j_{1},...,j_{k}}]-\operatorname{\mathbb{E}}[\mu_{j_{1},...,j_{k}}X_{i_{1},...,i_{k}}]-\operatorname{\mathbb{E}}[\mu_{i_{1},...,i_{k}}X_{j_{1},...,j_{k}}]-\mu_{i_{1},...,i_{k}}\mu_{j_{1},...,j_{k}}$
@@
Case 1: Pieces do not attack each other. First, we calculate how many such pairs there are. There are $n^{k}$ choices for $(i_{1},...,i_{k})$, which attack some $a_{i}n^{m}-b_{i}n^{k-c}$ spaces. So then, there are $n^{k}-a_{i}n^{m}+b_{i}n^{k-c}$ choices for $(j_{1},...,j_{k})$, meaning there are $n^{k}(n^{k}-a_{i}n^{m}+b_{i}n^{k-c})$ pairs that do not attack each other. Given that every piece attacks in some dimension less than $k$, the number of non-attacking pairs is always of degree $2k$, regardless of the dimension or type of piece.
A piece at $(i_{1},...,i_{k})$ attacks $a_{i}n^{m}-b_{i}n^{k-c}$ spaces, while a piece at $(j_{1},...,j_{k})$ attacks $a_{j}n^{m}-b_{j}n^{k-c}$ spaces. Then, the probability of a space being safe for a given $(i_{1},...,i_{k})$ and $(j_{1},...,j_{k})$ is ${\binom{n^{k}-a_{i}n^{m}-a_{j}n^{m}+b_{i}n^{k-c}+b_{j}n^{k-c}}{n^{k-c}}}/{\binom{n^{k}}{n^{k-c}}}$. This is the number of board configurations where pieces are placed at spaces that attack neither $(i_{1},...,i_{k})$ nor $(j_{1},...,j_{k})$, divided by all possible board configurations. Additionally, the probability of a bishop or queen being placed in the $r$th ring is $8r/n^{k}$. We evaluate the limit:
$\lim_{n\to\infty}\sum_{r_{k},r_{k}=0}^{(n-1)/2,(n-1)/2}...\sum_{r_{2},r_{2}=0}^{r_{3},r_{3}}\left(\frac{64r_{2}r_{2}}{n^{2k}}\prod_{\alpha=0}^{a_{i}n^{m}+a_{j}n^{m}-b_{i}n^{k-c}-b_{j}n^{k-c}}\frac{n^{k}-n-\alpha^{k-c}}{n^{k}-\alpha n^{k-c}}\right)$
$=\ \lim_{n\to\infty}\sum_{r_{k}=0}^{(n-1)/2}...\sum_{r_{2}=0}^{r_{3}}\left(\frac{8r_{2}}{n^{k}}\prod_{\alpha=0}^{a_{i}n^{m}-b_{i}n^{k-c}}\frac{n^{k}-n-\alpha}{n^{k}-\alpha}\right)\cdot$
$\sum_{r_{k}=0}^{(n-1)/2}...\sum_{r_{2}=0}^{r_{3}}\left(\frac{8r_{2}}{n^{k}}\prod_{\alpha=0}^{a_{j}n^{m}-b_{j}n^{k-c}}\frac{n^{k}-n-\alpha^{k-c}}{n^{k}-\alpha n^{k-c}}\right)$
@@
Case 2: Pieces attack each other. There are $n^{2k}-n^{k}$ distinct pairs of spaces. For pieces that attack $an^{m}$ spaces, we have $n^{2k}-an^{k+m}$ pairs that do not attack each other, meaning that there are $an^{k+m}-n^{k}$ pairs that do attack each other. However, as we take variance, we divide by $n^{2k}$, meaning that any terms of degree less than $n^{2k}$ tend to 0, so it is not necessary to calculate the percentage of safe spaces when two pieces attack each other. Therefore, we do not consider it as a term when we calculate variance.
Variance Conclusion We calculate $\operatorname{{Var}}(S_{n,k}/n^{k})=\operatorname{{Var}}(S_{n,k})/n^{2k}$. Then, from the previous sections:
@@
We note the constant terms cancel out, given that $\mu_{n}^{2}-\mu_{n}^{2}=0$. Then, all the other terms approach 0 as $n$ approaches infinity, so the variance can be made as close to 0 as desired, meaning the average percentage of safe spaces converges for $n^{k-1}$ line-pieces and $n$ hyper-pieces in any $k$ dimensions. ∎
@@
4. Line-Pieces in Higher Dimensions
As we consider higher dimensional chessboards, we first analyze pieces that attack linearly, placing $n^{k-1}$ pieces on a $k$-dimensional chessboard with side length $n$, as the number of pieces required to dominate the board is of order $n^{k-1}$.
4.1. Line-Rooks
Definition 4.1 (Line-Rooks).
For any $(i_{1},i_{2},\ldots,i_{k})$, there are knckn-c possible rook placements that attack it, noting that the constant $c$ term from overlap becomes negligible as $n$ approaches infinity. We place $n^{k-1}$ pieces and then divide by the total possible number of board combinations. We then use Lemma 2.1 to evaluate
Note that because of overlap, it does not take all $n^{k-1}$ line-rooks to dominate a $n^{k}$ board. As shown in [8], for $k=3$, only $n^{2}/2$ rooks are needed to dominate an $n\times n\times n$ chessboard, disregarding parity as $n$ tends to infinity. The number of rooks needed to dominate a general $k$-dimensional board remains one of many fascinatng open problems in the world of combinatorics and chess.
We can consider an alternate limit in $3$ dimensions, where instead of placing $n^{2}$ line-rooks, we place $n^{2}/2$ rooks. By Lemma 2.1, this becomes
4.2. Line-Bishops
Definition 4.2 (Line-Bishops in $k$ dimensions).
In $k$ dimensions, a $k$-dimensional line-bishop attacks as a normal bishop inside any plane it resides in, and does not attack any other spaces.
If we consider all planes that the line-bishop attacks on a $k$-dimensional chessboard, we find that we must take all combinations of $2$ dimensions out of the total $k$ dimensions. Hence, there are $(k^{2}-k)/2$ two-dimensional planes for a line-bishop to attack along. As we have defined it, a line-bishop placed in the center attacks as a regular bishop attackting 2n12n-1 spaces in each of these $(k^{2}-k)/2$ planes, for a total of $(k^{2}-k)n-(k^{2}-k)/2$ spaces attacked. Because $(k^{2}-k)/2$ is dwarfed as $n$ approaches infinity, we consider the central bishop as attacking $(k^{2}-k)n$ spaces.
$s\ :=\ 2\sum_{i=2}^{k}(i-1)r_{i}.$ (4.3)
As $n$ tends to infinity, we can apply Lemma 2.1 to see this is
4.3. Line-Queens
4.4. Variance of Line-Pieces
@@ 5. Hyper-Pieces in Higher Dimensions
Another way to extend the movement of chess pieces to $k$ dimensions is by exploring “hyper-pieces.” Any hyper-piece in $k$ dimensions attacks in (k1)(k-1)-dimensional subspaces. We place $n$ hyper-pieces in order to obtain dominance of the board. We are able to find the expected proportion of safe squares for hyper-rooks, but the other pieces evade our efforts beyond simple bounding arguments.
@@ 5.1. Hyper-rooks
Therefore, a hyper-rook attacks in a (k1)(k-1)-dimensional plane, which has $n^{k-1}$ spaces. Because the board has $k$ dimensions, there are $k$ of these planes, noting that we some $an^{k-2}$ for overlap points of all dimensions. This means that for $k$ dimensions, a hyper-rook leaves $n^{k}-kn^{k-1}-an^{k-2}$ safe spaces. Therefore, by Lemma 2.1, the expected proportion of safe spaces on a $k$-dimension board chessboard with side length $n$ and $n$ hyper-rooks as $n$ approaches infinity is
@@ 5.2. Hyper-bishops
Therefore, if we assume that all pieces see this maximal number of spaces, by Lemma 2.1, this gives us an expected percentage of safe spaces of

### v4

@@ Abstract.
The authors were supported by Emmanuel College Cambridge, The William & Mary Charles Center, The Cissy Patterson Fund, Williams College, The Finnerty Fund, and NSF Grant DMS2241623. Special thanks to Warren Johnson for helpful comments on an earlier draft and to the referees for many helpful comments.
@@ 1. Introduction
Miller et. al: Rooks in 2 Dimensions Section 2: Combinatorial Generalization Section 3: Bishops and Queens in 2 Dimensions Section 4: Hyper-pieces in Higher Dimensions Section 5: Line-pieces in Higher Dimensions
@@
and, using $\operatorname{\mathbb{E}}$ to denote the expected value function,
@@ 1.1. Notation and Basic Definitions in Higher Dimensions
We provide two natural extensions of chess pieces in higher dimensions, namely hyper-pieces and line-pieces. Hyper-pieces attack along (k1)(k-1)-dimensional hyperplanes; for example, a hyper-rook attacks any piece that shares one coordinate with it. Line-pieces attack along lines; for example, a line-rook attacks any piece that shares (k1)(k-1) coordinates with it.
We find that for hyper-rooks, after adjusting for their relative board covering power, they cover an equal amount of the board on average. More explicitly, we have the following theorem.
@@
The mean number of safe spaces on a $k$-dimensional chessboard with side length $n$ and $n$ hyper-rooks placed on it is asymptotically $n^{k}/e^{k}$. Hence, the probability a space is safe converges to $1/e^{k}$.
We then consider hyper-bishops and hyper-queens in $k$ dimensions. While a rook at any given space attacks the same number of spaces, the number of spaces a bishop attacks varies widely, often with little symmetry. Due to this, we are unable to determine the expected proportion of safe spaces in $k$ dimensions, and instead analyze $3$ dimensions.
Finally, when considering line-pieces, we determine that the expected proportion of safe squares when randomly placing $n^{k-1}$ line-rooks on a $k$-dimensional chessboard is $1/e^{k}$. We then consider line-bishops and line-queens and provide an integral for estimating the expected proportion of safe squares as $n$ approaches infinity. We computationally determine the expected proportion of safe squares when $k=3$. Our methods here are similar to the 2-dimensional case.
@@ 3.1. Probability of a Safe Square for Bishops and Queens
Interpreting this as a Riemann sum, with $x=r/n$, we can rewrite this as an integral, and then compute via integration by parts:
@@
$=8e^{-2}\left(\frac{e}{4}-\left(\frac{e}{4}-\frac{1}{4}\right)\right)$
$=\frac{2}{e^{2}}.$
Because we showed that constant terms do not impact the final limit, the expected proportion of safe spaces is
@@
We can define the even case similarly. The 0th ring has $4$ spaces, and the $r$th ring has $4(2r+1)$ spaces. Since we index from 0, the outermost ring is the (n/21)(n/2-1)st ring. A bishop in the 0th ring attacks n1n-1 spaces along one diagonal and n2n-2 along the other. For every ring movement outward, it attacks $2$ less spaces. Thus, the expected percentage of safe spaces on an even $n\times n$ chessboard with $n$ randomly placed bishops is
@@
Following the proof above, we observe that the limit of this expression as $n$ approaches infinity is $2/e^{4}$.
@@ 3.2. Variance for 2D Bishops and Queens
By the definition of variance and standard properties
@@
$=\ \frac{1}{n^{2k}}\left(\sum_{i_{1},i_{2},...,i_{k}=1}^{n}\operatorname{{Var}}(X_{i_{1},i_{2},...,i_{k}})+\sum_{\begin{subarray}{c}i_{1},...,i_{k},j_{1},...,j_{k}=1\\ (i_{1},...,i_{k})\neq(j_{1},...,j_{k})\end{subarray}}^{n}\operatorname{{Cov}}(X_{i_{1},i_{2},...,i_{k}},X_{j_{1},j_{2},...,j_{k}})\right).$ (3.8)
@@
Covariance Term We now consider the covariance term. $X_{i_{1},...,i_{k}}$ and $X_{j_{1},...,j_{k}}$ both consider all possible board configurations, so have the same expected value. Then the following holds:
@@
$=\ \operatorname{\mathbb{E}}[X_{i_{1},...,i_{k}}X_{j_{1},...,j_{k}}]-\operatorname{\mathbb{E}}[\mu_{j_{1},...,j_{k}}X_{i_{1},...,i_{k}}]-\operatorname{\mathbb{E}}[\mu_{i_{1},...,i_{k}}X_{j_{1},...,j_{k}}]+\mu_{i_{1},...,i_{k}}\mu_{j_{1},...,j_{k}}$
@@
Case 1: Pieces do not attack each other. First, we calculate how many such pairs there are. There are $n^{k}$ choices for $(i_{1},...,i_{k})$, which attack some $a_{i}n^{m}-b_{i}n^{k-c}$ spaces. Then there are $n^{k}-a_{i}n^{m}+b_{i}n^{k-c}$ choices for $(j_{1},...,j_{k})$, meaning there are $n^{k}(n^{k}-a_{i}n^{m}+b_{i}n^{k-c})$ pairs that do not attack each other. Given that every piece attacks in some dimension less than $k$, the number of non-attacking pairs is always of degree $2k$, regardless of the dimension or type of piece.
A piece at $(i_{1},...,i_{k})$ attacks $a_{i}n^{m}-b_{i}n^{k-c}$ spaces, while a piece at $(j_{1},...,j_{k})$ attacks $a_{j}n^{m}-b_{j}n^{k-c}$ spaces. Then the probability of a space being safe for a given $(i_{1},...,i_{k})$ and $(j_{1},...,j_{k})$ is ${\binom{n^{k}-a_{i}n^{m}-a_{j}n^{m}+b_{i}n^{k-c}+b_{j}n^{k-c}}{n^{k-c}}}/{\binom{n^{k}}{n^{k-c}}}$. This is the number of board configurations where pieces are placed at spaces that attack neither $(i_{1},...,i_{k})$ nor $(j_{1},...,j_{k})$, divided by all possible board configurations. Additionally, the probability of a bishop or queen being placed in the $r$th ring is $8r/n^{k}$. We evaluate the limit
$\lim_{n\to\infty}\sum_{r_{k},r_{k}=0}^{(n-1)/2,(n-1)/2}\ldots\sum_{r_{2},r_{2}=0}^{r_{3},r_{3}}\left(\frac{64r_{2}r_{2}}{n^{2k}}\prod_{\alpha=0}^{a_{i}n^{m}+a_{j}n^{m}-b_{i}n^{k-c}-b_{j}n^{k-c}}\frac{n^{k}-n-\alpha^{k-c}}{n^{k}-\alpha n^{k-c}}\right)$
$=\lim_{n\to\infty}\left[\sum_{r_{k}=0}^{(n-1)/2}\ldots\sum_{r_{2}=0}^{r_{3}}\left(\frac{8r_{2}}{n^{k}}\prod_{\alpha=0}^{a_{i}n^{m}-b_{i}n^{k-c}}\frac{n^{k}-n-\alpha}{n^{k}-\alpha}\right)\right]\cdot\left[\sum_{r_{k}=0}^{(n-1)/2}\ldots\sum_{r_{2}=0}^{r_{3}}\left(\frac{8r_{2}}{n^{k}}\prod_{\alpha=0}^{a_{j}n^{m}-b_{j}n^{k-c}}\frac{n^{k}-n-\alpha^{k-c}}{n^{k}-\alpha n^{k-c}}\right)\right]$
@@
Case 2: Pieces attack each other. There are $n^{2k}-n^{k}$ distinct pairs of spaces. For pieces that attack $an^{m}$ spaces, we have $n^{2k}-an^{k+m}$ pairs that do not attack each other, meaning that there are $an^{k+m}-n^{k}$ pairs that do attack each other. However, as we take the variance, we divide by $n^{2k}$, meaning that any terms of degree less than $n^{2k}$ tend to 0, so it is not necessary to calculate the percentage of safe spaces when two pieces attack each other. Therefore, we do not consider it as a term when we calculate the variance.
Variance Conclusion We calculate $\operatorname{{Var}}(S_{n,k}/n^{k})=\operatorname{{Var}}(S_{n,k})/n^{2k}$. Then, from the previous sections,
@@
We note the constant terms cancel out, given that $\mu_{n}^{2}-\mu_{n}^{2}=0$. All the other terms approach 0 as $n$ approaches infinity, so the variance can be made as close to 0 as desired, meaning the average percentage of safe spaces converges for $n^{k-1}$ line-pieces and $n$ hyper-pieces in any $k$ dimensions. ∎
@@ 4. Hyper-Pieces in Higher Dimensions
One way to extend the movement of chess pieces to $k$ dimensions is by exploring “hyper-pieces.” Any hyper-piece in $k$ dimensions attacks in (k1)(k-1)-dimensional subspaces. We place $n$ hyper-pieces in order to obtain dominance of the board. We are able to find the expected proportion of safe squares for hyper-rooks, but the other pieces evade our efforts beyond simple bounding arguments.
@@ 4.1. Hyper-rooks
Therefore, a hyper-rook attacks in a (k1)(k-1)-dimensional plane, which has $n^{k-1}$ spaces. Because the board has $k$ dimensions, there are $k$ of these planes, noting that we have counted some $an^{k-2}$ extra points from overlaps. This means that for $k$ dimensions, a hyper-rook leaves $n^{k}-kn^{k-1}-an^{k-2}$ safe spaces. Therefore, by Lemma 2.1, the expected proportion of safe spaces on a $k$-dimensional board with side length $n$ and $n$ hyper-rooks as $n$ approaches infinity is
@@ 4.2. Hyper-bishops
If we assume that all pieces see this maximal number of spaces, by Lemma 2.1, this gives us an expected percentage of safe spaces of
@@ 4.3. Hyper-queens
5. Line-Pieces in Higher Dimensions
It is also possible to consider pieces that move linearly in higher dimensions. We provide a broad definition of such pieces below.
Definition 5.1 (Higher dimensional line-pieces).
For any dimension $k>1$, a $k$-dimensional line-piece moves in perpendicular planes as it does on a $2$-dimensional chessboard. It can move in any $2$-dimensional plane, with all other coordinates being held constant. When determining the expected number of safe spaces, we place $n^{k-1}$ line-pieces.
Here we place $n^{k-1}$ line-pieces because the number of pieces required to dominate the board is of order $n^{k-1}$. If we were to place some $dn^{k-2}$ line-pieces, then each attacks some $a\cdot n$ spaces where $a$ is a constant factor, and when $n$ tends to infinity, $dan^{k-1}$ never covers a positive proportion of all $n^{k}$ spaces.
5.1. Line-Rooks
Definition 5.2 (Line-Rooks).
For any $(i_{1},i_{2},\ldots,i_{k})$, there are knckn-c possible rook placements that attack it, noting that the constant $c$ term from overlap becomes negligible as $n$ approaches infinity. We place $n^{k-1}$ pieces and then divide by the total possible number of board combinations. We use Lemma 2.1 to evaluate
Note that because of overlap, it does not take all $n^{k-1}$ line-rooks to dominate a $n^{k}$ board. As shown in [8], for $k=3$, only $n^{2}/2$ rooks are needed to dominate an $n\times n\times n$ chessboard, disregarding parity as $n$ tends to infinity. With this in mind, we can consider an alternate limit in $3$ dimensions, where instead of placing $n^{2}$ line-rooks, we place $n^{2}/2$ rooks. By Lemma 2.1, this becomes
The number of rooks needed to dominate a general $k$-dimensional board remains one of many fascinatng open problems in the world of combinatorics and chess.
5.2. Line-Bishops
Definition 5.3 (Line-Bishops in $k$ dimensions).
In $k$ dimensions, a $k$-dimensional line-bishop attacks as a 2-dimensional bishop inside any 2-dimensional plane it resides in, and does not attack any other spaces.
If we consider all planes that the line-bishop attacks on a $k$-dimensional chessboard, we find that we must take all combinations of $2$ dimensions out of the total $k$ dimensions. Hence, there are $(k^{2}-k)/2$ two-dimensional planes for a line-bishop to attack along. As we have defined it, a line-bishop placed in the center attacks as a regular bishop attacking 2n12n-1 spaces in each of these $(k^{2}-k)/2$ planes, for a total of $(k^{2}-k)n-(k^{2}-k)/2$ spaces attacked. Because $(k^{2}-k)/2$ is dwarfed as $n$ approaches infinity, we consider the central bishop as attacking $(k^{2}-k)n$ spaces.
$s\ :=\ 2\sum_{i=2}^{k}(i-1)r_{i}.$ (5.3)
As $n$ tends to infinity, we can apply Lemma 2.1 to see this is
5.3. Line-Queens
5.4. Variance of Line-Pieces